Find the magnitude and direction angle θ of the following vectors. Round the magnitude to the nearest tenth, and round the direction angle to the nearest degree, if rounding is necessary.
u = (9, -2)

Answer by stanbon(75887)

Find the magnitude and direction angle θ of the following vectors. Round the magnitude to the nearest tenth, and round the direction angle to the nearest degree, if rounding is necessary.
u = (9, -2)
----
magnitude = sqrt[9^2+2^2) = sqrt(85)
-------------------------------------
direction angle = arctan(-2/9) = 360-12.53 = 347.47 degrees
--------------
Cheers,
Stan H.
---------------